OpenStudy (anonymous):

Please! $15,000=$7,500(1+.0015/12)^12t... solve for "t"

OpenStudy (anonymous):

\[\large 15000=7500(1+0.0015)^{12t}\rightarrow 2=(1.0015)^{12t} \]

OpenStudy (anonymous):

so to solve for t:

OpenStudy (anonymous):

\[\large ln2=(12t)(ln1.0015) \] can you take it from here?

OpenStudy (anonymous):

what's "ln"?

OpenStudy (anonymous):

natural log...

OpenStudy (anonymous):

logarithm base e

OpenStudy (anonymous):

you really don't need ln... you can use common log.. log base 10

OpenStudy (anonymous):

I have not learned that yet... so I have no idea what to do after that :/

OpenStudy (anonymous):

how about common logs? log base 10?

OpenStudy (anonymous):

I don't know :(

OpenStudy (anonymous):

hmmm... that's strange that you're asked to solve for "t" without knowing about logarithms.

OpenStudy (anonymous):

it's an online class :/

OpenStudy (anonymous):

oh... so did the class go through logarithms?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it an online math class or online business class?

OpenStudy (anonymous):

Advanced Math with Financial Applications

OpenStudy (anonymous):

hmm... seeing those words advanced and financial makes me think they're assuming you already know about logarithms...

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

the only thing i can tell you here is that in order to solve for t in your question, you'll need to use logarithms...:(

OpenStudy (anonymous):

Well can you at least tell me how to input it in my calculator or something? or the link to a website that can teach about logarithms?

OpenStudy (anonymous):

sure...:) let's solve for t first on that last line I left you with...

OpenStudy (anonymous):

\[\large ln2=(12t)(ln1.0015)\] \[\large \frac{ln2}{12ln1.0015}=t\]

OpenStudy (anonymous):

ok... you have your calculator?

OpenStudy (anonymous):

do you see a "ln" button?

OpenStudy (anonymous):

or "log" button?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it doesn't matter which one you use... "ln" or "log"

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I don't think its the right answer considering t stands for time :/... .000086578

OpenStudy (anonymous):

no.... i got something else...

OpenStudy (anonymous):

does your calculator allow you to use parenthesis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good... make sure you enter: (ln2)/(12*ln1.0015)

OpenStudy (anonymous):

38.54?

OpenStudy (anonymous):

yep... that's what i got...:)

OpenStudy (anonymous):

YESSS! Thank you soo much!!!

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

you have a kweschun?

OpenStudy (anonymous):

You know how the original formula was : 15,000 = 7,500 (1 + .0015/ 12) ^12t... Well I'm writing it in my report and I saw we solved the logarithm of 1.0015 instead of 1.000125 which is the (1 + .0015/ 12) part :/

OpenStudy (anonymous):

\[\large 15000=7500(1+\frac{0.0015}{12})^{12t}\rightarrow 2=(1.0013)^{12t} \]

OpenStudy (anonymous):

\[\large 15000=7500(1+\frac{0.0015}{12})^{12t}\rightarrow 2=(1.00013)^{12t} \]

OpenStudy (anonymous):

Ok I did the same thing ... now I got 462 years? roughly?

OpenStudy (anonymous):

i got 444.354 years

OpenStudy (anonymous):

Ok I used 1.000125 not 1.00013 ... would that matter?

OpenStudy (anonymous):

haha...yours is more accurate...:)

OpenStudy (anonymous):

I'm ok with that haha, thanks again!

OpenStudy (anonymous):

yw...:)